OpenStudy (anonymous):

What's the derivative of 2xsinx and why?

OpenStudy (anonymous):

This problem is a combination of 4 formulas in derivatives of a function: 1. d [k f(x)] = k f'(x) dx 2. d (uv) = u dv + v du 3. d [x^n] = n x^(n-1) dx 4. d [sinx] = cos x dx

OpenStudy (anonymous):

we can use formula 1... where k=2 and f(x) = xsinx... d (2xsinx) = 2 d(xsinx) (1)

OpenStudy (anonymous):

we can use formula 2 in (1)... where u=x and v=sinx... 2 d(xsinx) = x d(sinx) dx + sinx d(x) dx (2)

OpenStudy (anonymous):

we can use formula 3 in u=x... that is du=(1)x^(1-1)=1 dx... we can use formula 4 in v=sinx... that is dv=cosx dx

OpenStudy (anonymous):

rewriting (1) combined with (2)... d (2xsinx) = 2 (xcosx + sinx) = 2xcosx + 2sinx... answer :)

OpenStudy (anonymous):

Ah! Thank you so much!

OpenStudy (anonymous):

you're welcome... :)